OpenStudy (anonymous):

She has 12 baskets left over from last year, and she plans to make 6 more each day. If there are 20 work days until the day she begins to sell the baskets, how many baskets will she have to sell?

OpenStudy (sodapop):

12 + 20 * 6 = 12 + 120 =132

OpenStudy (sodapop):

Hope that helped